# Nonlinear Schrödinger equation in terms of elliptic and hyperelliptic $\sigma$ functions

### 10 Mar 2024 v1

@@ 3.2. Sigma function and its derivatives
${\nu^{\mathrm{II}}_{j}}={\displaystyle\frac{1}{2y}}\sum_{k=j}^{2g-j}(k+1-j)\lambda_{k+1+j}x^{k}dx,\quad(j=1,\cdots,g)$
@@
$\theta\negthinspace\left[\begin{matrix}a\\ b\end{matrix}\right](z;\tau)=\sum_{n\in\mathbb{Z}^{g}}\exp\left[2\pi i\left\{{\displaystyle\frac{1}{2}}\ ^{t}\,\negthinspace(n+a)\tau(n+a)+\ ^{t}\,\negthinspace(n+a)(z+b)\right\}\right].$
The $\sigma$-function ([B1], p.336,[BEL1]), an analytic function on the space ${\mathbb{C}}^{g}$ and a theta series having modular invariance of a given weight with respect to $\mathfrak{M}$, is given by the formula
$\sigma(u)=\gamma_{0}\,\mathrm{exp}\left\{-{\displaystyle\frac{1}{2}}\ ^{t}u\eta^{\prime}{{\omega}^{\prime}}^{-1}u\right\}\theta\negthinspace\left[\begin{matrix}\delta^{\prime\prime}\\ \delta^{\prime}\end{matrix}\right]\left(\frac{1}{2}{{\omega}^{\prime}}^{-1}u\,;\,\mathbb{T}\right),$
where $\delta^{\prime}$ and $\delta^{\prime\prime}$ are half-integer characteristics giving the vector of Riemann constants with basepoint at $\infty$ and $\gamma_{0}$ is a certain non-zero constant. The $\sigma$-function vanishes only on $\kappa_{J}^{-1}(W_{X}^{g-1})$ (see for example [B1], p. 252). The hyperelliptic Weierstrass-$\wp$ and $\zeta$ functions are defined by
@@
As the Jacobi inversion formula,
@@
$\natural^{n}=\left\{\begin{array}[]{lll}n+1,n+3,\cdots,g-1&\text{for}\;g-n\equiv 0\;\mathrm{mod}\;2\\ n,n+2,\cdots,g&\text{for}\;g-n\equiv 1\;\mathrm{mod}\;2\\ \end{array}\right.$
@@
$\sigma_{\natural^{n}}=\bigg{(}\prod_{i\in\natural^{n}}\frac{\partial}{\partial u_{i}}\bigg{)}\sigma(u)$
@@
$\chi(\ell)$ $=\exp\big{\{}2\pi i\big{(}{}^{t}{\ell^{\prime}}\delta^{\prime\prime}-{}^{t}{\ell^{\prime\prime}}\delta^{\prime}+\tfrac{1}{2}{}^{t}{\ell^{\prime}}\ell^{\prime\prime}\big{)}\big{\}}\ (\in\{1,-1\})$
@@
(3.8) $\sigma_{\natural^{n}}(u+\ell)=\chi(\ell)\sigma_{\natural^{n}}(u)\exp L(u+\tfrac{1}{2}\ell,\ell)\ \ \hbox{for $u\in\kappa_{J}^{-1}(W_{X}^{n})$}.$
@@
(3.9) $\wp_{ij}(u+\ell)=\wp_{ij}(u),\quad\zeta_{i}(u+\ell)=\zeta_{i}(u)+\sum_{j=1}^{g}(\eta_{ij}^{\prime}\ell_{j}^{\prime}+\eta_{ij}^{\prime\prime}\ell_{j}^{\prime\prime}).$
@@ 4. The addition formulae
In this Section, we give the addition formulae of the hyperelliptic $\sigma$ functions which are the generalization of the addition formula (2.4). Those then give the key to construct the hyperelliptic solution of the Toda lattice.
4.1. The algebraic addition formula
Let us first recall the generalization of the addition formula (2.4):
@@ 4.1. The algebraic addition formula
$\begin{split}&\Psi_{n}((x_{1},y_{1}),\cdots,(x_{n},y_{n}))\\ &:=\left|\begin{matrix}1&\phi_{1}(x_{1},y_{1})&\cdots&\phi_{n-2}(x_{1},y_{1})&\phi_{n-1}(x_{1},y_{1})\\ 1&\phi_{1}(x_{2},y_{2})&\cdots&\phi_{n-2}(x_{2},y_{2})&\phi_{n-1}(x_{2},y_{2})\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 1&\phi_{1}(x_{n-1},y_{n-1})&\cdots&\phi_{n-2}(x_{n-1},y_{n-1})&\phi_{n-1}(x_{n-1},y_{n-1})\\ 1&\phi_{1}(x_{n},y_{n})&\cdots&\phi_{n-2}(x_{n},y_{n})&\phi_{n-1}(x_{n},y_{n})\\ \end{matrix}\right|,\end{split}$
@@
Let us first note that the curve $X$ has the algebraic addition structure [KMP, Ma6]:
Definition 4.2.
For given $P_{1},\cdots,P_{n}\in X$, we define
$\mu_{n}(P;P_{1},\cdots,P_{n})=\lim_{Q_{i}\to P_{i}}\frac{\Psi_{n+1}(P,Q_{1},\cdots,Q_{n})}{\Psi_{n}(Q_{1},\cdots,Q_{n})}$
for distinct $Q_{i}$’s.
Proposition 4.3.
For given $P_{1},\cdots,P_{n}\in X$, we have $Q_{i},\cdots,Q_{\ell}$ with $\ell=g$ for $n\geq g$ and $\ell=n$ otherwise, such that
$P_{1}+P_{2}+\cdots+P_{n}+Q_{1}+Q_{2}+\cdots+Q_{\ell}-(n+\ell)\infty\sim 0$
as a non trivial zero of $\mu_{n}(P;P_{1},\cdots,P_{n})$. For each $Q_{i}=(x_{i},y_{i})$, by letting $-Q_{i}=(x_{i},-y_{i})$, we have the addition property,
$P_{1}+P_{2}+\cdots+P_{n}-n\infty\sim(-Q_{1})+(-Q_{2})+\cdots+(-Q_{\ell})-(-\ell)\infty.$
Remark 4.4.
We should note that the Hyperelliptic involution $\iota:(x,y)\mapsto(x,-y)$, which induces $[-1]$-action on $J_{X}$ such that $u\mapsto-u$.
In this sense, since we fix the base point $\infty$, we use the notations, $n(x,y)$ for a point $(x,y)\equiv P\in X$, the addition $(x,y)+(x^{\prime},y^{\prime})$, and equality “$=$”; for example $(x,y)+(x^{\prime},y^{\prime})\neq(x+x^{\prime},y+y^{\prime})$ and $2(x,y)\neq(2x,2y)$ in general. Further we note that for every $P\in X$, $2P=-\sum_{i=1}^{2g-1}Q_{i}$ for points $Q_{i}$’s.
4.2. The analytic addition formula
@@
Assume that $(m,n)$ is a pair of positive integers. Let $(x_{i},y_{i})$ $(i=1,\cdots,m)$, $(x^{\prime}_{j},y^{\prime}_{j})$ $(j=1,\cdots,n)$ in $X$ and $u\in\kappa_{J}^{-1}(W_{X}^{m})$, $v\in\kappa_{J}^{-1}(W_{X}^{n})$ be points such that $u=w((x_{1},y_{1}),\cdots,(x_{m},y_{m}))$ and $v=w((x_{1}^{\prime},y_{1}^{\prime}),\cdots,(x_{n}^{\prime},y_{n}^{\prime}))$. Then the following relation holds :
(4.1) $\begin{split}&\frac{\sigma_{\natural^{m+n}}(u+v)\sigma_{\natural^{m+n}}(u-v)}{\sigma_{\natural^{m}}(u)^{2}\sigma_{\natural^{n}}(v)^{2}}\\ &=\delta(g,m,n)\frac{\prod_{i=0}^{1}\Psi_{m+n}((x_{1},y_{1}),\cdots,(x_{m},y_{m}),(x_{1}^{\prime},(-1)^{i}y_{1}^{\prime}),\cdots,(x_{n}^{\prime},(-1)^{i}y_{n}^{\prime}))}{(\Psi_{m}((x_{1},y_{1})\cdots,(x_{m},y_{m}))\Psi_{n}((x_{1}^{\prime},y_{1}^{\prime})\cdots,(x_{n}^{\prime},y_{n}^{\prime})))^{2}}\\ &\times\prod_{i=1}^{m}\prod_{j=1}^{n}\frac{1}{\Psi_{2}((x_{i},y_{i}),(x_{j}^{\prime},y^{\prime}_{j}))},\end{split}$
@@
Theorem 4.5 with $m=g$ and $n=1$ leads to the following Corollary:
Belokolos, Enolskii, and Salerno found the following relation [BES, Theorem 3.2];
@@
Let $(x_{i},y_{i})\in X$ $(i=1,\cdots,g)$, $({\widetilde{x}},{\widetilde{y}}):=(x^{\prime}_{1},y^{\prime}_{1})\in X$, $u\in{\mathbb{C}}^{g}$ and $v\in\kappa_{J}^{-1}(W_{X}^{1})$ such that $u=w((x_{1},y_{1}),\cdots,(x_{g},y_{g}))$ modulo $\Gamma_{X}$ and $v=w(({\widetilde{x}},{\widetilde{y}}))$ modulo $\Gamma_{X}$. Then we have
(4.2) $\begin{split}\frac{\sigma(u+v^{[1]})\sigma(u-v^{[1]})}{\sigma(u)^{2}\sigma_{\flat}(v^{[1]})^{2}}&={{\widetilde{x}}}^{g}-\sum_{i=1}^{g}\wp_{gj}(u){x^{\prime}_{1}}^{i+j-2}\\ &=F({\widetilde{x}})=({\widetilde{x}}-x_{1})({\widetilde{x}}-x_{2})\cdots({\widetilde{x}}-x_{g}).\end{split}$
Proof.
Direct computation shows it.
Noting the difference in the differential operators in Jacobian $J_{X}$, we generalize (2.5)-(2.9).
@@
    ${\sum_{i=1}^{g}e^{i-1}\left(\zeta_{i}(u+v)+\zeta_{i}(u-v)-2\zeta_{i}(u)\right)=-\sum_{\ell=1}^{g}\frac{2y_{\ell}F(e)}{(e-x_{\ell})(x-x_{\ell})F^{\prime}(x_{\ell})}=-F(e)\Delta_{x,e},}$
@@
where ${\zeta_{i}^{\natural_{j}}(v):=\left[\frac{\partial}{\partial u_{i}}\log\sigma_{\natural_{j}}(u)\right]_{u=v}}$, and $e$ is a generic parameter.
Proof.
Lemma 3.2 and (3.1) show 1. Noting $\partial/\partial x=\sum_{i}(2y/x^{i-1})\partial/\partial v_{i}$, we have 3, 4, and 5.
@@
Let $u={\widetilde{w}}(\iota_{X}(x_{1},y_{1}),\ldots,\iota_{X}(x_{g},y_{g}))$, and $v={\widetilde{w}}(\iota_{X}(x,y))$a.
(4.3) $\wp_{gg}(u+v)+\wp_{gg}(u-v)+2\wp_{gg}(u)=2\left[\sum_{r}\frac{y_{r}}{(x-x_{r})F^{\prime}(x_{r})}\right]^{2}+2\frac{f(x)}{F(x)^{2}}-2\lambda_{2g}-2x.$
Proof.
We consider the differential of 4.7 2 with respect to $u$. We interpret the differential (3.2) with respect to $u$ into differentials with respect to $x_{r}$, and then we have
$\begin{split}\wp_{gg}(u+v)&+\wp_{gg}(u-v)-2\wp_{gg}(u)=\sum_{\ell}\frac{2y_{\ell}}{F^{\prime}(x_{\ell})}\frac{\partial}{\partial x_{\ell}}\sum_{r}\frac{2y_{r}}{(x-x_{r})F^{\prime}(x_{r})}\\ &=2\left[\sum_{r}\frac{y_{r}}{(x-x_{r})F^{\prime}(x_{r})}\right]^{2}+\sum_{r}\frac{2}{F^{\prime}(x_{r})}\frac{\partial}{\partial x}\left[\frac{f(x)}{(x-x)F^{\prime}(x)}\right]\Bigr{|}_{x=x_{r}}\\ &=2\left[\sum_{r}\frac{y_{r}}{(x-x_{r})F^{\prime}(x_{r})}\right]^{2}+2\left[\frac{f(x)}{F(x)^{2}}-\lambda_{2g}-x-2\wp_{gg}(u)\right].\\ \end{split}$
@@
Proof.
The relation is proved by the integral.
@@
    The divisor of integrand in (4.4) is
    (4.5) $\left(\frac{f(x)}{(x-b_{r})F(x)^{2}}dx\right)=3\sum_{i=1,b_{i}}^{2g+1}(b_{i},0)-({\widetilde{x}},{\widetilde{y}})-({\widetilde{x}},-{\widetilde{y}})-2\sum_{i=1}^{g}(x_{i},y_{i})-2\sum_{i=1}^{g}(x_{i},-y_{i})-3\infty.$
@@
    By considering the local parameter $t$ ($t=x-{\widetilde{x}}$) at $({\widetilde{x}},\pm{\widetilde{y}})$, we have
@@
Proof.
The sketch of the proof is as follows. The right hand is the meromorphic function on $J_{X}\times{\widetilde{w}}(\iota_{X}X)$ because of the translational formula (3.9). Thus we show the equality of both sides. Let us show the right hand side is a meromorphic function of $S^{g}X\times X$. We note that both sides are odd function with respect to $v$. For $v$, we expand both sides,
$\zeta_{i}(u+v)-\zeta_{i}(u-v)=-2\sum_{ij}\wp_{ij}v_{j}-\frac{2}{3!}\sum_{ij}\wp_{ijk\ell}v_{j}v_{k}v_{\ell}+d_{>3}(v_{i}).$
The local parameter $t=x^{g+1}/y$ shows them.
@@
Lemma 4.11 is generalized as follows.
(4.7) $\begin{split}&[\zeta_{g-1}(u+v)-\zeta_{g-1}(u-v)+2{\widetilde{\zeta}}_{g-1}(v)]=\frac{2y}{F(x)}(x-\wp_{gg}),\\ &[\zeta_{g-2}(u+v)-\zeta_{g-2}(u-v)+2{\widetilde{\zeta}}_{g-2}(v)]=\frac{2y}{F(x)}(x^{2}-\wp_{gg}x+\wp_{g,g-1}).\end{split}$

### 14 Mar 2024 v2

@@ 3.2. Sigma function and its derivatives
${\nu^{\mathrm{II}}_{j}}=\frac{1}{2y}\sum_{k=j}^{2g-j}(k+1-j)\lambda_{k+1+j}x^{k}dx,\quad(j=1,\cdots,g)$
@@
$\theta\negthinspace\left[\begin{array}[]{c}a\\ b\end{array}\right](z;\tau)=\sum_{n\in\mathbb{Z}^{g}}\exp\left[2\pi i\left\{\frac{1}{2}\ ^{t}\,\negthinspace(n+a)\tau(n+a)+\ ^{t}\,\negthinspace(n+a)(z+b)\right\}\right].$
The $\sigma$-function ([B], p.336,[BEL1]), an analytic function on the space ${\mathbb{C}}^{g}$ and a theta series having modular invariance of a given weight with respect to $\mathfrak{M}$, is given by the formula
$\sigma(u)=\gamma_{0}\,\mathrm{exp}\left\{-\frac{1}{2}\ ^{t}u\eta^{\prime}{{\omega}^{\prime}}^{-1}u\right\}\theta\negthinspace\left[\begin{array}[]{c}\delta^{\prime\prime}\\ \delta^{\prime}\end{array}\right]\left(\frac{1}{2}{{\omega}^{\prime}}^{-1}u\,;\,\mathbb{T}\right),$
where $\delta^{\prime}$ and $\delta^{\prime\prime}$ are half-integer characteristics giving the vector of Riemann constants with basepoint at $\infty$ and $\gamma_{0}$ is a certain non-zero constant. The $\sigma$-function vanishes only on $\kappa_{J}^{-1}(W_{X}^{g-1})$ (see for example [B], p. 252). The hyperelliptic Weierstrass-$\wp$ and $\zeta$ functions are defined by
@@
As the Jacobi inversion formula, we have the relation,
@@
$\natural^{n}=\left\{\begin{array}[]{lll}n+1,n+3,\cdots,g-1&\mbox{for}\;g-n\equiv 0\;\mathrm{mod}\;2\\ n,n+2,\cdots,g&\mbox{for}\;g-n\equiv 1\;\mathrm{mod}\;2\\ \end{array}\right.$
@@
$\sigma_{\natural^{n}}=\bigg{(}\prod_{i\in\natural^{n}}\frac{\partial}{\partial u_{i}}\bigg{)}\sigma(u),$
@@
$\chi(\ell)$ $=$ $\exp\big{\{}2\pi i\big{(}{}^{t}{\ell^{\prime}}\delta^{\prime\prime}-{}^{t}{\ell^{\prime\prime}}\delta^{\prime}+\frac{1}{2}{}^{t}{\ell^{\prime}}\ell^{\prime\prime}\big{)}\big{\}}\ (\in\{1,-1\})$
@@
(3.5) $\sigma_{\natural^{n}}(u+\ell)=\chi(\ell)\sigma_{\natural^{n}}(u)\exp L(u+\frac{1}{2}\ell,\ell)\ \ \hbox{for $u\in\kappa_{J}^{-1}(W_{X}^{n})$}.$
@@
(3.6) $\wp_{ij}(u+\ell)=\wp_{ij}(u),\quad\zeta_{i}(u+\ell)=\zeta_{i}(u)+2\sum_{j=1}^{g}(\eta_{ij}^{\prime}\ell_{j}^{\prime}+\eta_{ij}^{\prime\prime}\ell_{j}^{\prime\prime}).$
@@ 4. The addition formulae
In this Section, we give the addition formulae of the hyperelliptic $\sigma$ functions which are the generalization of the addition formula (2.4). Those then provide the key to construct the hyperelliptic version of the algebraic differential relation (1.1) and (2.14).
Let us introduce the Frobenius-Stickelberger determinant:
@@
$\Psi_{n}((x_{1},y_{1}),\cdots,(x_{n},y_{n}))$
$:=\left|\begin{array}[]{cccccc}1&\phi_{1}(x_{1},y_{1})&\cdots&\phi_{n-2}(x_{1},y_{1})&\phi_{n-1}(x_{1},y_{1})\\ 1&\phi_{1}(x_{2},y_{2})&\cdots&\phi_{n-2}(x_{2},y_{2})&\phi_{n-1}(x_{2},y_{2})\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 1&\phi_{1}(x_{n-1},y_{n-1})&\cdots&\phi_{n-2}(x_{n-1},y_{n-1})&\phi_{n-1}(x_{n-1},y_{n-1})\\ 1&\phi_{1}(x_{n},y_{n})&\cdots&\phi_{n-2}(x_{n},y_{n})&\phi_{n-1}(x_{n},y_{n})\\ \end{array}\right|,$
@@
Assume that $(m,n)$ is a pair of positive integers. Let $(x_{i},y_{i})$ $(i=1,\cdots,m)$, $(x^{\prime}_{j},y^{\prime}_{j})$ $(j=1,\cdots,n)$ in $X$ and $u\in\kappa_{J}^{-1}(W_{X}^{m})$, $v\in\kappa_{J}^{-1}(W_{X}^{n})$ be points such that $u={\widetilde{w}}(\iota_{X}(x_{1},y_{1}),\cdots,\iota_{X}(x_{m},y_{m}))$ and $v={\widetilde{w}}(\iota_{X}(x_{1}^{\prime},y_{1}^{\prime}),\cdots,\iota_{X}(x_{n}^{\prime},y_{n}^{\prime}))$. Then the following relation holds :
$\frac{\sigma_{\natural^{m+n}}(u+v)\sigma_{\natural^{m+n}}(u-v)}{\sigma_{\natural^{m}}(u)^{2}\sigma_{\natural^{n}}(v)^{2}}$
$=\delta(g,m,n)\frac{\prod_{i=0}^{1}\Psi_{m+n}((x_{1},y_{1}),\cdots,(x_{m},y_{m}),(x_{1}^{\prime},(-1)^{i}y_{1}^{\prime}),\cdots,(x_{n}^{\prime},(-1)^{i}y_{n}^{\prime}))}{(\Psi_{m}((x_{1},y_{1})\cdots,(x_{m},y_{m}))\Psi_{n}((x_{1}^{\prime},y_{1}^{\prime})\cdots,(x_{n}^{\prime},y_{n}^{\prime})))^{2}}$
(4.2) ×$\times\prod_{i=1}^{m}\prod_{j=1}^{n}\frac{1}{\Psi_{2}((x_{i},y_{i}),(x_{j}^{\prime},y^{\prime}_{j}))},$
@@
Theorem 4.2 with $m=g$ and $n=1$ leads to the following Corollary, which Belokolos, Enolskii, and Salerno found [BES, Theorem 3.2];
@@
Let $(x_{i},y_{i})\in X$ $(i=1,\cdots,g)$, $(x,y):=(x^{\prime}_{1},y^{\prime}_{1})\in X$, $u\in{\mathbb{C}}^{g}$ and $v\in\kappa_{J}^{-1}(W_{X}^{1})$ such that $u=w((x_{1},y_{1}),\cdots,(x_{g},y_{g}))$ modulo $\Gamma_{X}$ and $v=w((x,y))$ modulo $\Gamma_{X}$. Then we have
(4.3) $\frac{\sigma(u+v)\sigma(u-v)}{\sigma(u)^{2}\sigma_{\natural^{1}}(v)^{2}}$ $=$ ${x}^{g}-\sum_{i=1}^{g}\wp_{gj}(u)x^{i-1}$
$=$ $F(x)=(x-x_{1})(x-x_{2})\cdots(x-x_{g}).$
Proof:  Direct computation shows it.
Considering the difference in the differential operators in Jacobian $J_{X}$, we generalize (2.5)-(2.9). Let $(x_{i},y_{i})\in X$ $(i=1,\cdots,g)$ and $(x,y)\in X$ as follows.
@@
    ${\sum_{i=1}^{g}e^{i-1}\left(\zeta_{i}(u+v)+\zeta_{i}(u-v)-2\zeta_{i}(u)\right)=-\sum_{\ell=1}^{g}\frac{2y_{\ell}F(e)}{(e-x_{\ell})(x-x_{\ell})F^{\prime}(x_{\ell})},}$
@@
where $e$ is a generic parameter. Further we let ${\zeta_{i}^{\natural_{j}}(v):=\left[\frac{\partial}{\partial{\widetilde{u}}_{i}}\log\sigma_{\natural^{j}}({\widetilde{u}})\right]_{u=v}}$ for ${\widetilde{u}}\in{\mathbb{C}}^{g}$.
Proof:  Lemma 3.2 and (3.1) show 1. Noting $\partial/\partial x=\sum_{i}(2y/x^{i-1})\partial/\partial v_{i}$, we have 3, 4, and 5.
@@
Let $u={\widetilde{w}}(\iota_{X}(x_{1},y_{1}),\ldots,\iota_{X}(x_{g},y_{g}))$, and $v={\widetilde{w}}(\iota_{X}(x,y))$.
(4.4) $\wp_{gg}(u+v)$ $+$ $\wp_{gg}(u-v)+2\wp_{gg}(u)=$
22 22 $\left[\sum_{r}\frac{y_{r}}{(x-x_{r})F^{\prime}(x_{r})}\right]^{2}+2\frac{f(x)}{F(x)^{2}}-2\lambda_{2g}-2x.$
Proof:  We consider the differential of Lemma 4.4 (ii) with respect to $u$. We interpret the differential (3.2) with respect to $u$ into differentials with respect to $x_{r}$, then we have
$\wp_{gg}(u+v)$ $+$ $\wp_{gg}(u-v)-2\wp_{gg}(u)=\sum_{\ell}\frac{2y_{\ell}}{F^{\prime}(x_{\ell})}\frac{\partial}{\partial x_{\ell}}\sum_{r}\frac{2y_{r}}{(x-x_{r})F^{\prime}(x_{r})}$
$=$ $2\left[\sum_{r}\frac{y_{r}}{(x-x_{r})F^{\prime}(x_{r})}\right]^{2}+\sum_{r}\frac{2}{F^{\prime}(x_{r})}\frac{\partial}{\partial x}\left[\frac{f(x)}{(x-x)F^{\prime}(x)}\right]\Bigr{|}_{x=x_{r}}$
$=$ $2\left[\sum_{r}\frac{y_{r}}{(x-x_{r})F^{\prime}(x_{r})}\right]^{2}+2\left[\frac{f(x)}{F(x)^{2}}-\lambda_{2g}-x-2\wp_{gg}(u)\right].$
@@
Proof:  The relation is proved by the integral.
@@
    The divisor of the integrand in (4.5) is
    (4.6) $\left(\frac{f(x)}{(x-b_{r})F(x)^{2}}dx\right)$ $=$ $3\sum_{i=1,b_{i}}^{2g+1}(b_{i},0)-({\widetilde{x}},{\widetilde{y}})-({\widetilde{x}},-{\widetilde{y}})$
    -- $2\sum_{i=1}^{g}(x_{i},y_{i})-2\sum_{i=1}^{g}(x_{i},-y_{i})-3\infty.$
@@
    By considering the local parameter $t$ ($t=x-{\widetilde{x}}$) at $({\widetilde{x}},\pm{\widetilde{y}})$, we obtain
@@
Proof:  The left hand side is the meromorphic function on $J_{X}\times W_{X}^{1}$ because of the translational formula (3.6). The right hand side is a meromorphic function on $S^{g}X\times X$ and thus, on $J_{X}\times W_{X}^{1}$ by the Abel-Jacobi theorem. Thus we show the equality of both sides. We compare the divisors. As a function of $v$, both sides vanish at every finite branch point $(x,y)=(b_{r},0)$ from the translational formula (3.6) and the definition of ${\widetilde{\zeta}}_{g}$. At $v=0$ or $P=\infty$, both sides diverge in order of one. At $v=u$ or $v=-u$, correspoing to $x=x_{i}$, the order of the pole is one. We note that both sides are odd functions with respect to $v$. For $v$, we expand both sides at $v=0$,
$\zeta_{g}(u+v)-\zeta_{g}(u-v)=-2\sum_{j}\wp_{gj}v_{j}-\frac{2}{3!}\sum_{jk}\wp_{gjk\ell}v_{j}v_{k}v_{\ell}+\mbox{the oder of }v\geq 3$
The local parameter $t=x^{g+1}/y$ shows
$\zeta_{g}(u+v)-\zeta_{g}(u-v)-2{\widetilde{\zeta}}_{g}(v)=2(x_{1}+\cdots+x_{g})t+\frac{2}{t}+\mbox{the oder of }t>0.$
$\frac{2{\widetilde{y}}}{F({\widetilde{x}})}=\frac{2}{t}\left(1+(x_{1}+\cdots+x_{g})t^{2}+\mbox{the oder of }t>0\right).$
The leading terms of both sides match. Hence we prove it.
@@
Lemma 4.8 may be generalized as follows.
$[\zeta_{g-1}(u+v)-\zeta_{g-1}(u-v)+2{\widetilde{\zeta}}_{g-1}(v)]=\frac{2y}{F(x)}(x-\wp_{gg}),$
(4.8) $[\zeta_{g-2}(u+v)-\zeta_{g-2}(u-v)+2{\widetilde{\zeta}}_{g-2}(v)]=\frac{2y}{F(x)}(x^{2}-\wp_{gg}x+\wp_{g,g-1}).$